OpenStudy (anonymous):

write √-4+10 in a+bi form

OpenStudy (anonymous):

Do you understand how the square root of a negative number works?

OpenStudy (anonymous):

- Nope .

OpenStudy (anonymous):

The square root of negative 1 is defined as the imaginary number "i". I'm sure that sounds weird at first, but don't worry :) \[\sqrt{-1} = i\] And... \[(\sqrt{-1})^{2} = i^{2}\] which means that...\[(\sqrt{-1})^{2} = -1 = i^{2}\]

OpenStudy (anonymous):

So, any time you have a square root with a negative number under it, just imagine it as the positive number multiplied by a negative 1. For example:\[\sqrt{-25} = \sqrt{(25)*(-1)}\] And then you can take the square roots of each part... \[\sqrt{-25} = \sqrt{(25)*(-1)} = \sqrt{25}*\sqrt{-1} = 5 * i\]

OpenStudy (anonymous):

- oh okay , i get that .

OpenStudy (anonymous):

Good :) So, your question is: Write √-4+10 in a+bi form So, using what we just talked about, what is the square root of -4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\sqrt{-4} = \sqrt{(4)(-1)} = (\sqrt{4})(\sqrt{-1})\]

OpenStudy (anonymous):

- Okay .

OpenStudy (anonymous):

Square root of -4 is square root of 4 multiplied by square root of negative 1. The square root of 4 is 2 (you got that part right!) and square root of negative 1 is "i". So:\[\sqrt{-4} = \sqrt{(4)(-1)}=\sqrt{4}\sqrt{-1} = 2i\]

OpenStudy (anonymous):

- Okay .

OpenStudy (anonymous):

To write √-4 + 10 in "a + bi" form, you just write the real number part (10, in this problem) as the "a" and the imaginary part (2i in this problem) as the "bi" part. so: √-4 + 10 = 10 + 2i

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

- iget it .

OpenStudy (anonymous):

Great :) Glad to help...